x^2+y^2=1/4 and (x+1/2)^2+y^2+z^2=1...Then how does one get from second equation -1

x^2+y^2=1/4   and   (x+1/2)^2+y^2+z^2=1...Then how does one get from second equation -1<=z<=1..

Grade:

1 Answers

Askiitians_Expert Yagyadutt
13 years ago

hello darshana

 

your question is not clear...please make it more clear what to find outSmile

 

(x+1/2)^2+y^2+z^2=1   this is equation of a sphere...with limitation on z as given  by   -1 <= z <=1  

and  x^2 + y^2 = 1/4 is equation of a circle ...

 

now what to do ?

 

i will post the solution after u make it clear okk !

Regards

 

yagya

 

askiitians_expert